OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (anonymous):

i really can do this one bc i have a caulator but not a scientific caulator so i cant do the fraction part can some one help

OpenStudy (sleepyjess):

Divide both sides by -3 x+7 = 4 so then you have x+7 = 4 x+7 = -4 x = -3 x = -11

OpenStudy (sleepyjess):

I skipped actually showing one of the steps To get the x on it's own, you have to subtract 7 from both sides

OpenStudy (sleepyjess):

But wait... you already posted this question and it got answered in Linear Algebra

OpenStudy (anonymous):

ik but my computer froz and shut off so i re poseted it so idk